OpenStudy (czarluc):

{log3(x)+2^(y+2)=1 {log3(x^2)+2^y=3 Solve the system

OpenStudy (alekos):

well I suppose you can start by subtracting as the equations stand

OpenStudy (welshfella):

yes that would eliminate 2^y

OpenStudy (welshfella):

are those logs to the base 3?

OpenStudy (czarluc):

yes

OpenStudy (alekos):

then solve for x after some log rule manipulation

OpenStudy (czarluc):

so the equation would be -log3(x) +2^2=-2?

OpenStudy (alekos):

I'll give you the first step and then I want you to try the second

OpenStudy (czarluc):

okay

OpenStudy (czarluc):

the term 2^(y+2) is a correction

OpenStudy (czarluc):

the "+2" is part of the exponent of 2

OpenStudy (alekos):

stand by

OpenStudy (alekos):

OK bottom equation becomes 2logx + 2^y = 3 and now we multiply the top equation by 2 and then subtract the equations to eliminate 2logx

OpenStudy (czarluc):

then the equation becomes 4^(y+2)-2^y=-1

OpenStudy (alekos):

No that's way out, but I just tried to solve for y and I get the log of a -ve number which is undefined, so that can only mean you've given me the wrong equations. I reckon the 1 and 3 need to be swapped

OpenStudy (czarluc):

Its the exact equations given, maybe it's just undefined

OpenStudy (alekos):

If that's the case then there is no solution. Have they given you an answer?

OpenStudy (czarluc):

No its only an example. the system is just inconsistent

OpenStudy (alekos):

Yeah, that's another way of putting it